OpenStudy (anonymous):

a fraction reduces to 36 if its numerator is (6x)5 , what is its denominator?

OpenStudy (anonymous):

30x

OpenStudy (anonymous):

@kumar2006 that wouldn't be true that would be 30/30= 1 not 36

OpenStudy (anonymous):

It would be 5/6 x

OpenStudy (lgbasallote):

do you mean the numerator is (6x)^5? lez say a is your denominator...unknow (6x)^5/a = 36 (6x)^5 = 36a (6^5 x^5)/6^2 = a 6^3 x^5 = a that's your denominator

OpenStudy (anonymous):

30x/(5/6 x) = 36